OpenStudy (anonymous):

what exactly is a 'real number' does that mean whole # or can it be like .5?

OpenStudy (lstrasmann):

yeah, i think real numbers can be basically any number on the number line (including decimals and fractions) according to wikipedia it includes integers fractions and unrational numbers (Pi for example)

OpenStudy (anonymous):

whole # - this is not defined by mathematicians (use integer, positive integer, natural number, etc)

OpenStudy (turingtest):

a real number is any number that appears on the real line; i.e. does not have an imaginary (complex) part complex numbers involve \[i=\sqrt{-1}\]which is not a real number, because there is no real number \((-\infty,\infty)\) number that satisfies \(\sqrt{-1}\) our guesses might be 1 and -1, but...\[1\cdot1=1\neq-1\]and\[(-1)(-1)=1\neq-1\]so there is no "real" number that satisfies the expression real numbers in no way have to be whole numbers

OpenStudy (anonymous):

so, .5 is a real number?

OpenStudy (turingtest):

does it have any complex part in it ? an \(i\) or \(\sqrt1\) ?

OpenStudy (turingtest):

I mean \(\sqrt{-1}\)

OpenStudy (anonymous):

0.5 + 0i

OpenStudy (turingtest):

oh you cheater :P

OpenStudy (anonymous):

:-)

OpenStudy (turingtest):

it must have a non-zero complex part to be exact

OpenStudy (turingtest):

...if it is *not* to be a real number, that is

OpenStudy (anonymous):

But isn't the real line part of the complex plane?

OpenStudy (turingtest):

all real numbers are complex, not all complex are real if I am phrasing it poorly above I am only being too lazy since I'm about to leave, sorry :/

OpenStudy (anonymous):

s'K, just kidding a bit, it's the algebraist in me......

OpenStudy (turingtest):

good habit ;)

OpenStudy (anonymous):

I bet OP wish she never asked now.....:-)

OpenStudy (anonymous):

so is .5 a real number???

OpenStudy (anonymous):

A simple answer is that pretty much all the numbers you are used to (including 0.5) are real.